# ShieldAgent: Shielding Agents via Verifiable Safety Policy Reasoning

### v1

@@ Appendix H Prompt Template
[…]
Extraction Guidelines: Do not summarize, modify, or simplify any part of the original policy. Copy the exact descriptions. Ensure each extracted policy is self-contained and can be fully interpreted by looking at its Definition, Scope, and Policy Description. If the Definition or Scope is unclear, leave the value as None. Avoid grouping multiple policies into one block. Extract policies as individual pieces of statements.
Provide the output in the following JSON format: ```json
@@
[…]
Predicate Formatting: Each predicate must include: Predicate Name: Use snake_case format. Description: A brief, clear explanation of what the predicate represents. Keywords: A list of descriptive keywords providing relevant context (e.g., actions, entities, attributes).
LTL Symbol Definitions: Always: ALWAYS Eventually: EVENTUALLY Next: NEXT Until: UNTIL Not: NOT And: AND Or: OR Implies: IMPLIES
Output Format: ```json
@@
Output Requirements: Ensure each rule is explicitly defined and unambiguous. Keep predicates general when applicable (e.g., use create_project instead of click_create_project). Avoid combining unrelated rules into a single LTL statement.
[…]
Output Requirements: Provide step-by-step reasoning under the section Reasoning. Include the label on whether the predicate is good, needs refinement, or redundant. If refinement is needed, provide a structured JSON including: Updated predicate with definitions and keywords. Each of the updated rules which are associated with the updated predicate. Definitions of the predicate in each rule’s context.
Output Format: Reasoning: 1. Step-by-step reasoning for why the predicate is good, needs refinement, or is redundant. 2. If yes, then reason about how to refine or remove the redundant predicate. Decision: Yes/No If yes, then provide the following: Output JSON:
@@
[…]
Output Requirements: Provide step-by-step reasoning under the section Reasoning. Include a decision label on whether the predicates should be merged. If merging is needed, provide a structured JSON including: Updated predicates with definitions and keywords. Updated rules with the new merged predicates.
Output Format: Reasoning: 1. Step-by-step reasoning for why the predicates should or should not be merged. 2. If merging is needed, explain how the predicates and rules were updated to ensure completeness and consistency. Decision: Yes/No If yes, then provide the following: Output JSON:

### v2

@@ Appendix G Case Study
Figure 7: Illustration of an unsafe trajectory and its corresponding annotation from a sample in our ShieldAgent-Bench. Each sample contains one safe trajectory and two unsafe trajectories generated by two distinct attack algorithms. As shown on the left, each trajectory comprises multiple steps of agent-environment interaction, where the environment observation includes both textual information (e.g., HTML AX-tree and terminal stdout) and the agent’s action consists of a CoT reasoning process followed by an executable action. As shown on the right, we carefully annotate each agent action step with a task success label, safety violation label, and corresponding risk categories.
@@
Figure 11: An example of the verifiability refinement stage of our safety policy structure optimization algorithm, where a compound rule (process_content) is decomposed into multiple atomic rules that are more concrete and verifiable (publish_content, download_content, create_content). Specifically, the decomposition process takes into account the broader context of the original rule, including its NLP descriptions and document source, to ensure accuracy and fidelity.
Figure 12: An example of the verifiability refinement stage in our safety policy structure optimization algorithm, where the original extracted rule which contains a vague predicate (comply_with_laws) is refined into a more specific and grounded rule with an updated predicate (comply_with_GDPR_laws). With the optimized predicate, the refined rule could explicitly guide the agent to invoke a relevant GDPR checking tool during the guardrail process, enabling a more accurate verification result.
Figure 13: An example of the redundancy pruning stage in our safety policy structure optimization algorithm, where two clustered rules containing predicates with identical contextual implications but different names (publish_personal_data and disclose_personal_data) are merged such that they share a single predicate (publish_personal_data). This pruning operation reduces redundancy in the rule space and improves the efficiency of the verification process.
Figure 14: An end-to-end example of the guardrail procedure in ShieldAgent. Given the task instruction, interaction history, and invoked action as input, (1) ShieldAgent first extracts relevant action predicates from the agent’s output and matches all higher-level action nodes within the action tree. (2) It then retrieves relevant rule circuits for all identified actions and (3) searches for similar verification workflows from the long-term memory module. Based on this, (4) ShieldAgent generates a step-by-step shielding plan to assign predicate values by invoking operations from a comprehensive tool library (illustrated in detail in Fig. 15, Fig. 16, and Fig. 17). Once all predicates are assigned, (5) it then generates shielding code (shown in Fig. 18) to formally verify each rule. Finally, (6) it performs probabilistic inference to determine the safety label and (7) reports all violated rules along with corresponding explanations and remediation suggestions (detailed in Fig. 19).
Figure 15: An example illustrating the shielding plan generation process of ShieldAgent. This example focuses on guardrailing the unsafe trajectory from the dataset sample shown in Fig. 7. Specifically, given the retrieved action rule circuits corresponding to the extracted actions across different hierarchical levels (matched via a static action tree), ShieldAgent first performs relevance analysis to filter out irrelevant rules and resolve ambiguities within the rule predicates. It then invokes tools to construct a predicate truth table, listing all predicate names and definitions for assignments.
Figure 16: An example (following Fig. 15) illustrating the shielding plan generation process of ShieldAgent. In this stage, ShieldAgent first assigns predicates that can be directly inferred from the context (e.g., target action predicates such as update_bio). Then for more sophisticated state predicates (e.g., data_is_harmful), ShieldAgent invokes various verification tools following the MCP protocol to determine and assign their truth values.
Figure 17: An example (following Fig. 16) illustrating the shielding plan generation process of ShieldAgent. In this stage, ShieldAgent first updates the predicate truth table with the newly assigned predicates. Then, it proceeds to verify state predicates where no specialized tools are available by leveraging generic query-based verification tools to generate individual queries for each predicate and assign their values based on the responses. Afterward, ShieldAgent invokes an additional tool to verify whether all predicates have been successfully assigned or to generate a new plan for any remaining unassigned predicates.
Figure 18: An example (following Fig. 17) illustrating the shielding code generation process of ShieldAgent. Once all predicates in the truth table have been assigned, ShieldAgent then generates model checking code for each rule (e.g., based on prover9) and transports these codes via corresponding tools to formally verify them. It then collects the verification results along with detailed explanations and remediation feedback for any violated rules.
Figure 19: An example (following Fig. 18) illustrating the safety certification process of ShieldAgent. After all rules have been verified, ShieldAgent performs safety certification to estimate the safety probability of executing the invoked action. It then determines a safety label by comparing this probability against a predefined threshold using a prescribed certification method (e.g., barrier function). Finally, ShieldAgent reports the safety label along with any violated rules, accompanied by detailed explanations and remediation suggestions.
@@ Appendix H Prompt Template
Prompt Template for Policy Extraction system: You are a helpful policy extraction model to identify actionable policies from organizational safety guidelines. Your task is to exhaust all the potential policies from the provided organization handbook which sets restrictions or guidelines for user or entity behaviors in this organization. You will extract specific elements from the given guidelines to produce structured and actionable outputs. user: As a policy extraction model to clean up policies from {organization (e.g. GitLab)}, your tasks are: 1.: Read and analyze the provided safety policies carefully, section by section. 2.: Exhaust all actionable policies that are concrete and explicitly constrain behaviors. 3.: For each policy, extract the following four elements: 1.: Definition: Any term definitions, boundaries, or interpretative descriptions for the policy to ensure it can be interpreted without any ambiguity. These definitions should be organized in a list. 2.: Scope: Conditions under which this policy is enforceable (e.g. time period, user group). 3.: Policy Description: The exact description of the policy detailing the restriction or guideline. 4.: Reference: All the referenced sources in the original policy article from which the policy elements were extracted. These sources should be organized piece by piece in a list. Extraction Guidelines: –: Do not summarize, modify, or simplify any part of the original policy. Copy the exact descriptions. –: Ensure each extracted policy is self-contained and can be fully interpreted by looking at its Definition, Scope, and Policy Description. –: If the Definition or Scope is unclear, leave the value as None. –: Avoid grouping multiple policies into one block. Extract policies as individual pieces of statements. Provide the output in the following JSON format: ```json
@@
Prompt Template for Linear Temporal Rule Extraction system: You are an advanced policy translation model designed to convert organizational policies into structured Linear Temporal Logic (LTL) rules. Your task is to extract verifiable rules from the provided safety guidelines and express them in a machine-interpretable format while maintaining full compliance with logical correctness. user: As a policy-to-LTL conversion model, your tasks are: 1.: Carefully analyze the policy’s definition, scope, and policy description. 2.: Break down the policy into structured rules that precisely capture its constraints and requirements. 3.: Translate each rule into LTL using atomic predicates derived from the policy. Translation Guidelines: –: Use atomic predicates that are directly verifiable from the agent’s observations and action history. –: Prefer positive predicates over negative ones (e.g., use store_data instead of is_data_stored). –: If a rule involves multiple predicates, decompose it into smaller, verifiable atomic rules whenever possible. –: Emphasize action-based predicates, ensuring that constrained actions are positioned appropriately within logical expressions (e.g., “only authorized users can access personal data” should be expressed as: $(\texttt{is\_authorized}\wedge\texttt{has\_legitimate\_need})\Rightarrow\texttt{access\_personal\_data}$ (8) ). Predicate Formatting: Each predicate must include: –: Predicate Name: Use snake_case format. –: Description: A brief, clear explanation of what the predicate represents. –: Keywords: A list of descriptive keywords providing relevant context (e.g., actions, entities, attributes). LTL Symbol Definitions: –: Always: ALWAYS –: Eventually: EVENTUALLY –: Next: NEXT –: Until: UNTIL –: Not: NOT –: And: AND –: Or: OR –: Implies: IMPLIES Output Format: ```json
@@
Output Requirements: –: Ensure each rule is explicitly defined and unambiguous. –: Keep predicates general when applicable (e.g., use create_project instead of click_create_project). –: Avoid combining unrelated rules into a single LTL statement.
Prompt Template for Verifiability Refinement (VR) system: You are a helpful predicate refinement model tasked with ensuring predicates in the corresponding rules are clean, verifiable, concrete, and accurate enough to represent the safety policies. Your task is to verify each predicate and refine or remove it if necessary. user: As a predicate refinement model, your tasks are: 1. Check if the provided predicate satisfies the following criteria: –: Verifiable: It should be directly verifiable from the agent’s observation or action history. –: Concrete: It should be specific and unambiguous. –: Accurate: It must represent the intended fact or condition precisely. –: Atomic: It should describe only one fact or action. If it combines multiple facts, break it into smaller predicates. –: Necessary: The predicate must refer to meaningful information. If it is redundant or assumed by default, remove it. –: Unambiguous: If the same predicate name is used in different rules but has different meanings, rename it for clarity. 2. If refinement is needed, refine the predicate accordingly with one of the following: –: Rewrite the predicate if it is unclear or inaccurate. –: Break it down into smaller atomic predicates if it combines multiple facts or conditions. –: Rename the predicate to reflect its context if it is ambiguous. –: Remove the predicate if it is redundant or unnecessary for the rule. Output Requirements: –: Provide step-by-step reasoning under the section Reasoning. –: Include the label on whether the predicate is good, needs refinement, or redundant. –: If refinement is needed, provide a structured JSON including: *: Updated predicate with definitions and keywords. *: Each of the updated rules which are associated with the updated predicate. *: Definitions of the predicate in each rule’s context. Output Format: Reasoning: 1. Step-by-step reasoning for why the predicate is good, needs refinement, or is redundant. 2. If yes, then reason about how to refine or remove the redundant predicate. Decision: Yes/No If yes, then provide the following: Output JSON:
@@
Prompt Template for Redundancy Pruning (RP) system: You are a helpful predicate merging model tasked with analyzing a collection of similar predicates and their associated rules to identify whether there are at least predicates that can be merged or pruned. Your goal is to simplify and unify rule representation while ensuring the meaning and completeness of the rules remain intact after modifying the predicates. user: As a predicate merging model, your tasks are: 1.: Identify predicates in the cluster that can be merged based on the following conditions: –: Redundant Predicates: If two or more predicates describe the same action or condition but use different names or phrasing, merge them into one. –: Identical Rule Semantics: If two rules describe the same behavior or restriction but are phrased differently, unify the predicates and merge their logics to represent them with fewer rules. 2.: Ensure the merged predicates satisfy the following: –: Consistency: The merged predicate must be meaningful and represent the combined intent of the original predicates. –: Completeness: The new rules must perfectly preserve the logic and intent of all original rules. Output Requirements: –: Provide step-by-step reasoning under the section Reasoning. –: Include a decision label on whether the predicates should be merged. –: If merging is needed, provide a structured JSON including: *: Updated predicates with definitions and keywords. *: Updated rules with the new merged predicates. Output Format: Reasoning: 1.: Step-by-step reasoning for why the predicates should or should not be merged. 2.: If merging is needed, explain how the predicates and rules were updated to ensure completeness and consistency. Decision: Yes/No If yes, then provide the following: Output JSON: